# NeurIT: Pushing the Limit of Neural Inertial Tracking for Indoor Robotic IoT

### v1

@@ Abstract
Inertial tracking is vital for robotic IoT and has gained popularity thanks to the ubiquity of low-cost Inertial Measurement Units (IMUs) and deep learning-powered tracking algorithms. Existing works, however, have not fully utilized IMU measurements, particularly magnetometers, nor maximized the potential of deep learning to achieve the desired accuracy. To enhance the tracking accuracy for indoor robotic applications, we introduce NeurIT, a sequence-to-sequence framework that elevates tracking accuracy to a new level. NeurIT employs a Time-Frequency Block-recurrent Transformer (TF-BRT) at its core, combining the power of recurrent neural network (RNN) and Transformer to learn representative features in both time and frequency domains. To fully utilize IMU information, we strategically employ body-frame differentiation of the magnetometer, which considerably reduces the tracking error. NeurIT is implemented on a customized robotic platform and evaluated in various indoor environments. Experimental results demonstrate that NeurIT achieves a mere 1-meter tracking error over a 300-meter distance. Notably, it significantly outperforms state-of-the-art baselines by 48.21% on unseen data. NeurIT also performs comparably to the visual-inertial approach (Tango Phone) in vision-favored conditions and surpasses it in plain environments. We believe NeurIT takes an important step forward toward practical neural inertial tracking for ubiquitous and scalable tracking of robotic things. NeurIT, including the source code and the dataset, is open-sourced here: https://github.com/NeurIT-Project/NeurIT.
@@
neural inertial tracking, indoor robotic localization, inertial measurement unit, time-frequency learning.
@@ I Introduction
Indoor tracking is vital for robots, as it allows them to provide location-based services and navigate in areas where satellite signals are weak or unavailable. Numerous techniques have been investigated for this purpose. Vision-based methods [1] show remarkable performance in favorable environments, yet degrade in plain environments and bad lighting conditions.
@@
LiDAR has been used to generate accurate tracking and mapping [2], which, however, is too costly and may suffer from severe drift in the absence of environmental feature points. Wireless tracking solutions utilizing WiFi [3], millimeter-wave [4], or ultra-wide bandwidth (UWB) [5] technologies have also been widely studied, yet the performance is not robust in real-world cases while the deployment overhead is significant. These methodologies either rely on external infrastructure or depend on environmental features. In contrast, inertial tracking using IMUs has carved a niche by offering a truly infrastructure-free solution [6]. An IMU is a self-contained module integrating accelerometer, gyroscope, and magnetometer, which is a standard component in ubiquitous devices including commercial robots. IMU-based tracking method does not depend on external infrastructure, rendering an inside-out solution that is suitable for indoor robotic navigation in various conditions.
@@
However, achieving high accuracy in inertial tracking is challenging due to sensor noise and drifts. Conventional studies examine filtering techniques like Kalman filter [7], complementary filter [8], etc., to reduce IMU noise. These techniques may not work well in distinct conditions or for long duration. The emergence of machine learning, specifically neural inertial tracking based on deep neural networks, offers a new approach that promises improved precision and generalizability.
Two types of neural networks are mainly exploited: RNN and Transformers. Despite acceptable results, they face limitations, particularly on cross-domain or unseen datasets, leaving considerable scope for refinement to realize the full potential of deep learning in inertial tracking.
@@
  • RNN-based solutions can capture temporal features well [9, 10, 11, 12]. However, their operational inefficiency and the short-term memory causing vanishing gradients [13] can destabilize learning outcomes.
@@
  • Transformers for inertial tracking [14, 15] have shown promising results. However, they struggle to capture information beyond the perception field, especially essential past data for the current motion context. Additionally, performance drops on unseen datasets indicate insufficient generalization capabilities based on their reports.
@@
  • Existing works mostly employ accelerations and gyroscopes only and exclude the magnetometer for learning. Few works have integrated magnetometers for drift compensation, yet merely employ complementary filter-based techniques [16, 12], yielding only marginal enhancements in the outcomes.
@@
To overcome these limitations, we present NeurIT, a novel system that unleashes the full potential of neural inertial tracking and achieves remarkable accuracy for robotic tracking. NeurIT excels mainly in two aspects: First, based on an in-depth understanding of inertial data properties, NeurIT incorporates magnetometer data as an additional feature with accelerometer and gyroscope measurements. While it has been believed that the magnetometer is susceptible to indoor interference [17, 18, 19] and is excluded from learning previously, we notice that it has a lower drift than the gyroscope in providing long-term orientation information, rendering it a valuable input for potential drift compensation. It is noticeable that magnetometers primarily capture pose information, while accelerations and gyroscopes depict the movement at a fine-grained level. To align these features together, we integrate magnetometer differentiation within the body frame. This ensemble enhances the network’s ability to counteract the noise-induced drift in pose estimation. Second, to maximize the potential of neural networks for enhanced inertial tracking, we propose a novel Time-frequency Block-Recurrent Transformer, named TF-BRT, which embraces the advantages of both RNN and Transformer and overcomes their respective drawbacks by extracting pre-sequence information and local motion information within the perception field. Particularly, TF-BRT extracts both time and frequency domain information to obtain a more comprehensive understanding of object movement. By learning from complete IMU data with an advanced neural network design, NeurIT achieves robust tracking capabilities that endure extended duration and cover substantial distances.
@@
To implement and evaluate NeurIT, we build a customized robotic platform, which consists of a commodity 9-axis IMU for NeurIT, a LiDAR-inertial system for ground truth, and a visual-inertial system for comparison. We verify the performance of NeurIT on two public benchmarks and a newly built dataset. Our own dataset is collected in three different campus buildings, totaling 33.7 km over about 15 hours. The experimental results demonstrate that NeurIT not only outperforms all baseline models in the existing benchmark dataset, it also achieves better results under more challenging, practical conditions. Specifically, compared with the best baseline, NeurIT improves the tracking accuracy by 48.21% on unseen data and achieves an average drift rate of 0.62%, indicating its superior robustness over long distances. Additionally, NeurIT achieves the thus far best performance on two benchmark datasets, RIDI [20] and RONIN [9]. Further, when compared to commercial visual-inertial algorithm [21], NeurIT exceeds in plain environments while achieving comparable performance in visual-favored environments. Fig. 1 illustrates the superior tracking accuracy achieved by NeurIT, while previous neural inertial tracking algorithm (IONet [10]) and traditional control algorithm (EKF [7]) fail to provide robust tracking outcomes. This enhances the potential of inertial tracking as a viable solution for various indoor tracking applications and invites new perspectives on the longstanding challenge of indoor positioning.
@@
    We propose NeurIT, an inertial tracking system that leverages a full set of IMU data under a novel network design named TF-BRT. TF-BRT exploits the advantages of both RNN and Transformer and effectively learns from time-frequency domain information, enabling accurate tracking in various environments and minimizing drift over extended periods and distances.
@@
    We implement NeurIT with a customized robotic platform and test it in various environments using a newly built dataset. NeurIT surpasses the existing baselines on all the benchmark datasets and performs well against a visual-inertial tracking system. We have open-sourced the collected dataset and source code for the research community.
In the rest of the paper, we investigate the insights of neural inertial tracking in §II and present the system design in §III. We present implementation details in §IV, followed by experiment settings in §V. We evaluate NeurIT scrupulously in §VI. Finally, we review the literature in §VII and conclude the paper in §VIII.
@@
Figure 2: Global magnetometers study in the indoor environment. The first row shows the heading error comparison between global magnetometers and gyroscopes, which indicates that magnetometers can be used for orientation drift compensation from gyroscopes. The second row illustrates the orientation of global magnetometers on the floor plane.
II Observations and Measurements
@@
IMU-based tracking often suffers from orientation drift, which is mainly caused by gyroscope noises. In outdoor scenarios, global magnetometers can correct the drift to some extent. Nevertheless, in indoor environments, ferromagnetic materials in building structures can distort magnetometers [17, 18]. This potential distortion serves as a potential consideration, possibly explaining the omission of magnetometers as input features in prior neural inertial tracking studies. To test this hypothesis, we collected magnetometer and gyroscope data from different indoor environments and used them to estimate the heading orientations. We then compared the accuracy of each method by calculating the heading error over time and plotting the results in Fig. 2. The first row represents the heading error outcomes given by global magnetometers and gyroscopes. From the dashed lines, despite the impact of complex buildings on global magnetometers and the presence of fluctuating orientation errors, the long-term orientation drift of global magnetometers is minimal compared to gyroscope-derived heading directions. This intriguing finding leads us to posit that magnetometers could serve as a viable compensatory feature in the context of neural inertial tracking.
@@
Nevertheless, how to encapsulate the three modalities remains a challenge, given their unique perspectives on measuring the target’s motion. First, for precise floor plane tracking, converting accelerations and gyroscopes from the IMU body frame to the Earth frame is crucial since the IMU operates in a non-inertial coordinate system. Unlike accelerometers and gyroscopes, which measure the amount of motion but not the precise direction, magnetometers can directly infer the orientation of an object. To improve the alignment of magnetometers with accelerometers and gyroscopes, we introduce the derivative of magnetometers in the body frame, which encompasses both the pose information and the object’s orientation motion, enhancing the network’s ability to learn motion patterns at a finer-grained level. We believe it can be a powerful compensation feature for indoor tracking tasks, culminating in the sensor-fusion algorithm for our neural inertial tracking method. In section §VI-E, we conduct a study with magnetometers, which shows the superior performance of incorporating magnetometer data into our system.
@@
III NeurIT Design
III-A Overview
@@
NeurIT presents a sequence-to-sequence framework to reconstruct the moving trajectories of robots using low-cost IMU sensors only, without relying on any external infrastructure (e.g., Access Points, cellular stations). Fig. 3 overviews the design of NeurIT. To track the robot’s movement, IMU readings that consist of accelerations, gyroscopes, and magnetometers are collected and pre-processed into a time series representation. The core of NeurIT is a neural network called TF-BRT, a novel model that leverages block-recurrent attention [22] and time-frequency learning. The block-recurrent attention mechanism allows the neural network to incorporate both the current IMU data and the historical information to produce a more accurate tracking result. Moreover, the input includes both time and frequency domain features, enhancing its inference ability for different motion states. At the end of NeurIT, the output generated by TF-BRT, which is a sequence of movement speed and orientation, is integrated to calculate the robot’s location at any time. We consider predicting the robot’s location directly is not an effective option because locations do not constitute a finite set. Additionally, we introduce a multi-loss learning method and data augmentation to further enhance the robustness of TF-BRT.
III-B Data Preparation
@@
IMU coordinate frame, which changes all the time, reports the raw IMU data. The Earth coordinate frame, on the other hand, defines the moving trajectory. To resolve this discrepancy, coordinate transformation is required. A common practice in neural inertial tracking [9] is to transform both the raw IMU data and the ground truth into the same coordinate frame, which is called the navigation frame. We use the same coordinate frame transformation as RONIN [9] in NeurIT, which uses a heading-agnostic coordinate frame (Earth coordinate frame) to represent the IMU data.
The coordinate transformation is only applied to accelerations and gyroscopes:
$(\textbf{a}^{g},\textbf{w}^{g})=\textbf{q}(\textbf{a}^{b},\textbf{w}^{b})\textbf{q}^{*},$ (1)
@@
where a and w denote the accelerations and gyroscopes, and $b$ and $g$ represent the body frame and Earth coordinate frame. In contrast, we retain the magnetometers in the body frame of IMU and apply a differential operation:
$\dot{\mathbb{m}}^{b}$ =˙$=\dot{\mathbb{R}}\cdot\mathbb{m}^{g}$ (2)
$=\mathbb{\Omega}\cdot\mathbb{R}\cdot\mathbb{m}^{g}$
$=\begin{pmatrix}0&-w_{z}&w_{y}\\ w_{z}&0&-w_{x}\\ -w_{y}&w_{x}&0\end{pmatrix}\cdot\mathbb{R}\cdot\mathbb{m}^{g},$
@@
where $\mathbb{\Omega}$ is the angular velocity matrix. Obviously, with the hidden rotation information, $\dot{\mathbb{m}}^{b}$ can be used to compensate for gyroscope drift as described in the aforementioned context.
@@
Following with the combination of three-modality data as a time series, $\textbf{X}=[\textbf{a}^{g};\textbf{w}^{g};\dot{\textbf{m}}^{b}]$, which is then segmented into sequences of length $L$ through a sliding window. The operation is done before the input, generating data samples $\textbf{X}_{n}\in\mathbb{R}^{L\times D},n=[1,2,3,\cdots]$, where $D$ is the IMU data dimension.
III-C Model Architecture

### v2

@@ Abstract
Inertial tracking is vital for robotic IoT and has gained popularity thanks to the ubiquity of low-cost inertial measurement units and deep learning-powered tracking algorithms. Existing works, however, have not fully utilized IMU measurements, particularly magnetometers, nor have they maximized the potential of deep learning to achieve the desired accuracy. To address these limitations, we introduce NeurIT, which elevates tracking accuracy to a new level. NeurIT employs a Time-Frequency Block-recurrent Transformer (TF-BRT) at its core, combining both RNN and Transformer to learn representative features in both time and frequency domains. To fully utilize IMU information, we strategically employ body-frame differentiation of magnetometers, considerably reducing the tracking error. We implement NeurIT on a customized robotic platform and conduct evaluation in various indoor environments. Experimental results demonstrate that NeurIT achieves a mere 1-meter tracking error over a 300-meter distance. Notably, it significantly outperforms state-of-the-art baselines by 48.21% on unseen data. Moreover, NeurIT demonstrates robustness in large urban complexes and performs comparably to the visual-inertial approach (Tango Phone) in vision-favored conditions while surpassing it in feature-sparse settings. We believe NeurIT takes an important step forward toward practical neural inertial tracking for ubiquitous and scalable tracking of robotic things. NeurIT is open-sourced here: https://github.com/aiot-lab/NeurIT.
@@
neural inertial tracking, indoor localization, inertial measurement unit, time-frequency learning.
@@ I Introduction
Indoor tracking is vital for robots, as it allows them to provide location-based services and navigate in areas where satellite signals are weak or unavailable. Numerous techniques have been investigated for this purpose. Vision-based methods [1] perform well in good conditions but struggle in simple environments and poor lighting.
@@
LiDAR has been used to generate accurate tracking and mapping [2], which, however, is too costly and may drift without feature points. Wireless tracking solutions utilizing WiFi [3], millimeter-wave [4], etc, have also been widely studied, yet the performance is not robust in real-world cases, while the deployment overhead is significant. These methodologies either rely on external infrastructure or depend on environmental features. In contrast, inertial tracking using inertial measurement units (IMUs) has carved a niche by offering a truly infrastructure-free solution [5, 6]. An IMU is a self-contained module integrating accelerometer, gyroscope, and magnetometer, which is a standard component in ubiquitous devices, including commercial robots. The IMU-based tracking method does not depend on external infrastructure, rendering an inside-out solution that is suitable for indoor robotic navigation in various conditions.
@@
However, achieving highly accurate inertial tracking is challenging due to sensor noise and drift. Conventional studies examine filtering techniques like Kalman filter [7], complementary filter [8], etc, to reduce IMU noise. These techniques may not work well in distinct conditions or for long durations. The emergence of machine learning, specifically neural inertial tracking based on deep neural networks, offers a new approach that promises improved precision and generalizability.
Two types of neural networks are mainly exploited: recurrent neural networks (RNNs) and Transformers. Despite acceptable results, their performance drops in cross-domain or unseen scenarios, indicating room for further improvement.
@@
  • RNN-based solutions can capture temporal features well [9, 10, 11, 12]. However, they suffer from limited memory and operational inefficiencies, with issues such as vanishing gradients [13] often destabilizing the learning process.
@@
  • Transformer-based approaches [14, 15] have shown promising results. Nevertheless, they struggle to capture information beyond the perception field, especially essential past data for the current motion context.
@@
  • Existing works mostly employ accelerations and gyroscopes only, and exclude the magnetometer for learning. Few works have integrated magnetometers for drift compensation, yet merely employ complementary filter-based techniques [16, 12], yielding only marginal enhancements in the outcomes.
@@
To overcome these limitations, we present NeurIT, a robust system that unleashes the full potential of neural inertial tracking and achieves remarkable accuracy for robotic tracking. NeurIT excels mainly in two aspects: First, based on an in-depth understanding of inertial data properties, NeurIT incorporates magnetometer data as an additional feature with accelerometer and gyroscope measurements. Although the magnetometer has previously been believed to be susceptible to indoor interference [17, 18, 19] and was excluded from learning, we noticed that it has a lower drift than the gyroscope in providing long-term orientation information, making it a valuable input for drift compensation. It is noticeable that magnetometers primarily capture pose information, while accelerations and gyroscopes depict the movement at a fine-grained level. To align these features together, we integrate magnetometer differentiation within the body frame. This ensemble enhances the network’s ability to counteract the noise-induced drift in pose estimation. Second, to maximize the potential of neural networks for enhanced inertial tracking, we propose Time-frequency Block-Recurrent Transformer, named TF-BRT, which embraces the advantages of both RNN and Transformer and overcomes their respective drawbacks by extracting pre-sequence information and local motion information within the perception field. Particularly, TF-BRT extracts both time and frequency domain information to obtain a more comprehensive understanding of object movement. By learning from complete IMU data with an advanced neural network design, NeurIT achieves robust tracking capabilities that endure extended duration and cover substantial distances.
@@
To implement and evaluate NeurIT, we build a customized robotic platform, which consists of a commodity 9-axis IMU for NeurIT, a LiDAR-inertial system for ground truth, and a visual-inertial system for comparison. We verify the performance of NeurIT on two public benchmarks and a newly built dataset. Our dataset is collected in three different campus buildings, totaling 33.7 km over 15 hours. The experimental results demonstrate that NeurIT not only outperforms all baseline models in the existing benchmark dataset, but it also achieves better results under more challenging conditions. Specifically, compared to the best baseline, NeurIT improves the tracking accuracy by 48.21% on unseen data and achieves an average drift rate of 0.62%, indicating its superior robustness over long distances. Additionally, NeurIT achieves the thus far best performance on two benchmark datasets, RIDI [20] and RONIN [9]. To evaluate its real-world applicability, we test NeurIT in two large urban complexes, a shopping mall and a concert hall, where it maintains strong robustness in complex environments. Further, when compared to a commercial visual-inertial system [21], NeurIT exceeds in plain environments while achieving comparable performance in visual-favored environments. Fig. 1 illustrates the superior tracking accuracy achieved by NeurIT, while prior neural inertial tracking approach (IONet [10]) and traditional control method (EKF [7]) fail to provide robust tracking outcomes. This enhances the potential of inertial tracking as a viable solution for various indoor tracking applications, inviting new perspectives on the longstanding challenge of indoor positioning.
@@
    We propose NeurIT, an inertial tracking system that fully utilizes IMU data under an advanced network named TF-BRT. TF-BRT combines the strengths of RNNs and Transformers to effectively learn time-frequency features, enabling accurate tracking across diverse environments while minimizing long-term drift.
@@
    We implement NeurIT on a customized robotic platform and test it in various environments using a newly built dataset. NeurIT surpasses the existing baselines on all the benchmarks and performs well against a visual-inertial tracking system. We have open-sourced the collected dataset and source code for the research community.
In the rest of the paper, we review the literature in §II. Then we investigate the insights of neural inertial tracking in §III and present the system design in §IV. Implementation details are given in §V, followed by experiment settings in §VI. We evaluate NeurIT scrupulously in §VII. Finally, we conclude the paper in §VIII.
@@
Figure 2: Global magnetometers study. First row: heading error comparison between global magnetometers and gyroscopes; Second row: visualization of global magnetometers.
II Related Works
Neural Inertial Tracking: IONet [10] first integrates neural networks for inertial tracking, which greatly improves the accuracy compared with traditional methods [16, 12]. With its success, RONIN [9] and DeepVIP [22], etc, continue to explore the limit of RNN for inertial tracking and achieve promising results. After that, CTIN [14], A2DIO [15], etc, replicate the success of the Transformer in the NLP field and further improve the tracking accuracy. Additionally, researchers have developed models optimized for mobile deployment [23], while LLIO [24] has been specifically designed for wheeled vehicles. To the best of our knowledge, NeurIT is the first work to combine the advantages of RNN and Transformer for neural inertial tracking.
Magnetometers Usage: Most aforementioned neural inertial tracking works use accelerations and gyroscopes exclusively since magnetometers will be affected by indoor environments [18]. Some works, like $A^{3}$ [25] and MUSE [16], utilize magnetometers to calibrate gyroscope drift. In addition, Tlio [26] incorporates deep learning to enhance the robustness of EKF. However, the fundamental idea of these works is under the framework of filtering techniques, which is distinct from NeurIT. NeurIT utilizes differentiation of body-frame magnetometers to achieve fusion and complementarity of IMU data from a neural network input perspective.
Fused Inertial Tracking: Despite the impressive capabilities of modern neural networks, accumulating path errors over time remains a challenge for precise indoor positioning. To address this, integrating indoor maps is essential for achieving high-precision localization. For instance, NILoc [27] employs a transformer-based architecture to determine device location from velocity sequences, while [28] uses spatial map embeddings with IMU data for indoor location prediction. Other approaches, such as combining IMU estimations with WiFi [29, 30] or acoustic signals [31], offer limited improvements and are affected by ambient interference. Visual [32, 33] and LiDAR-based [34] tracking solutions have shown promise in both indoor and outdoor environments. However, these methods are prone to drift and tracking failure in feature-sparse areas. To enhance system robustness, researchers have incorporated IMU information, resulting in visual-inertial [35] and LiDAR-inertial [2] tracking systems.
III Observations and Measurements
@@
IMU-based tracking often suffers from orientation drift, primarily caused by gyroscope noise. While global magnetometers can help correct this drift outdoors, indoor environments pose challenges due to magnetic distortions from ferromagnetic building materials [17, 18]. This distortion may explain why prior neural inertial tracking studies often exclude magnetometer inputs. To test this hypothesis, we collected magnetometer and gyroscope data from different indoor environments to estimate the heading orientations. We compared their accuracy by calculating the heading error over time, as shown in Fig. 2. The first row presents heading errors from global magnetometers and gyroscopes. Despite fluctuations and distortions caused by complex indoor structures, the dashed lines reveal that magnetometers exhibit significantly less long-term drift than gyroscopes. This intriguing finding suggests that magnetometers may serve as a valuable complementary feature for neural inertial tracking.
@@
Nevertheless, effectively integrating the three modalities remains challenging due to their distinct perspectives on motion sensing. First, for precise floor plane tracking, converting accelerations and gyroscopes from the IMU body frame to the Earth frame is crucial since the IMU operates in a non-inertial coordinate system. Unlike accelerometers and gyroscopes, which measure the amount of motion but not the precise direction, magnetometers can directly infer the orientation of an object. To improve the alignment of magnetometers with accelerometers and gyroscopes, we introduce the derivative of magnetometers in the body frame, which encompasses both the pose information and the object’s orientation motion, enhancing the network’s ability to learn motion patterns at a finer-grained level. In §VII-E, we present a study demonstrating that integrating magnetometer data significantly enhances the performance of our system.
@@
IV NeurIT Design
IV-A Overview
@@
NeurIT presents a sequence-to-sequence framework to reconstruct the moving trajectories of robots using IMU sensors only, without relying on any external infrastructure (e.g., access points, cellular stations). Fig. 3 overviews the design of NeurIT. To track the robot’s movement, IMU readings that consist of accelerations, gyroscopes, and magnetometers are collected and pre-processed into a time series representation. The core of NeurIT is a neural network called TF-BRT, which leverages block-recurrent attention [36] and time-frequency learning. The block-recurrent attention mechanism enables the model to leverage both current IMU data and historical context for more accurate tracking. By incorporating features from both time and frequency domains, the model improves its ability to handle diverse motion states. At the end of NeurIT, TF-BRT outputs a sequence of velocities and orientations, which are integrated to compute the robot’s location over time. We avoid direct location prediction, as location is a continuous variable without a finite set. Additionally, we introduce multi-loss learning and data augmentation to further enhance the robustness of TF-BRT.
IV-B Data Preparation
@@
The IMU coordinate frame, which constantly changes, reports the raw IMU data. The Earth coordinate frame, on the other hand, defines the moving trajectory. To resolve this discrepancy, a coordinate transformation is required. A common practice in neural inertial tracking is to transform the raw IMU data and the ground truth into the same coordinate frame called the navigation frame [9]. We use the same coordinate frame transformation as RONIN [9] in NeurIT, which uses a heading-agnostic coordinate frame (Earth coordinate frame) to represent the IMU data.
@@
The coordinate transformation is only applied to accelerations and gyroscopes: $(\textbf{a}^{g},\textbf{w}^{g})=\textbf{q}(\textbf{a}^{b},\textbf{w}^{b})\textbf{q}^{*}$, where a and w denote the accelerations and gyroscopes, $b$ and $g$ represent the body frame and Earth coordinate frame. q is the quaternion, and $\textbf{q}^{*}$ is its conjugate, with both provided by the IMU. Conversely, we keep the magnetometers in the body frame and apply differentiation operation ($\dot{\Box}$):
$\dot{\mathbf{m}}^{b}$ $=\dot{\mathbf{R}}\cdot\mathbf{m}^{g}$ (1)
$=\mathbf{\Omega}\cdot\mathbf{R}\cdot\mathbf{m}^{g}$
$=\begin{pmatrix}0&-w_{z}&w_{y}\\ w_{z}&0&-w_{x}\\ -w_{y}&w_{x}&0\end{pmatrix}\cdot\mathbf{R}\cdot\mathbf{m}^{g},$
@@
where $\dot{\mathbb{m}}^{b}$ is the time derivative of the magnetometer vector in the body frame, $\mathbb{R}$ is the rotation matrix from the Earth frame to the body frame, and $\mathbb{\Omega}$ is the angular velocity matrix formed from the gyroscope readings. Obviously, with the hidden rotation information, $\dot{\mathbf{m}}^{b}$ can be used to compensate for gyroscope drift as described in the aforementioned context.
@@
The combined three-modality time-series data, denoted as $\textbf{X}=[\textbf{a}^{g};\textbf{w}^{g};\dot{\textbf{m}}^{b}]$, is segmented into fixed-length sequences of length $L$ using a sliding window. These segments constitute the data samples for the model input $\textbf{X}_{n}\in\mathbb{R}^{L\times D},n=[1,2,3,\cdots]$, where $D$ is the IMU data dimension.
IV-C Model Architecture